If the measures of the angles of a triangle are in the ratio of 1:2:3, and if the perimeter of the triangle is 30+10√3, what is the length of the smallest side?
Answer by stanbon(75887)

If the measures of the angles of a triangle are in the ratio of 1:2:3, and if the perimeter of the triangle is 30+10√3, what is the length of the smallest side?
Angles:: x + 2x + 3x = 180
6x = 180
x = 30 degrees (smallest angle)
2x = 60 degrees and 3x = 90 degrees
------------
So the sides are in the ratio of 1:2:Sqrt(3)
----
Equation:
x + 2x + sqrt(3)x = 30+10sqrt(3)
3x + sqrt(3)x = 30+10sqrt(3)
---
So, x = 10 (That is the length of the smallest side)
-----------
Cheers,
Stan H.
---------